# A Survey on Efficient Inference for Large Language Models

### v2

@@ 2.3 Efficiency Analysis
(Sec. 6.1), snode [Speculative Decoding, snode] [Graph and Operator Optimization, snode] ] ] [Model-level Optimization (Sec. 5), mnode_1 [Model Compression
@@ 5.2.1 Quantization
In weight-only quantization, GPTQ [190] represents an early advancement in LLM quantization, building upon the traditional algorithm OBQ [215]. OBQ utilizes an optimal quantization order per row of the weight matrix, guided by the reconstruction error relative to the Hessian matrix of unquantized weights. After each quantization step, OBQ iteratively adjusts the unquantized weights to mitigate reconstruction errors. However, the frequent updating of the Hessian matrix during quantization escalates computational complexity. GPTQ streamlines this process by adopting a uniform left-to-right order for quantizing each row, thus circumventing the need for extensive Hessian matrix updates. This strategy substantially reduces computational demands by computing the Hessian matrix solely during the quantization of one row, then leveraging the computing results for subsequent rows, expediting the overall quantization procedure. LUT-GEMM [191] presents a novel dequantization method utilizing a Look-Up Table (LUT), aiming to accelerate the inference process of quantized LLMs by reducing the dequantization overhead. Additionally, it adopts a non-uniform quantization approach known as Binary-Coding Quantization (BCQ), which incorporates learnable quantization intervals. AWQ [192] observes that weight channels vary in importance for performance, particularly emphasizing those aligned with input channels exhibiting outliers in activations. To enhance the preservation of critical weight channels, AWQ utilizes a reparameterization method. This technique selects reparameterization coefficients via grid search to minimize reconstruction errors effectively. OWQ [193] observes the difficulty of quantizing weights associated with activation outliers. To address this challenge, OWQ employs a mixed-precision quantization strategy. This method identifies weak columns in the weight matrix and allocates higher precision to these specific weights, while quantizing the rest of the weights at a lower precision level. SpQR [194] introduces a methodology where weight outliers are identified and allocated higher precision during quantization, while the rest of the weights are quantized to 3 bits. SqueezeLLM [195] proposes to store the outliers in a full-precision sparse matrix, and apply non-uniform quantization to the remaining weights. The values for non-uniform quantization are determined based on quantization sensitivity, which contributes to improved performance of the quantized model. QuIP [196] introduces LDLQ, an optimal adaptive method for a quadratic proxy objective. The study reveals that ensuring incoherence between weight and Hessian matrices can enhance the effectiveness of LDLQ. QuIP utilizes LDLQ and achieves incoherence by employing random orthogonal matrix multiplication. FineQuant [197] utilizes a heuristic approach to determine the granularity of quantization per column, combining empirical insights gained from experiments to design a quantization scheme. QuantEase [198] builds upon GPTQ. When quantizing each layer, it proposes a method based on Coordinate Descent to compensate for the unquantized weights more precisely. Additionally, QuantEase can leverage quantized weights from GPTQ as an initialization and further refine the compensation process. LLM-MQ [199] protects the weight outliers with FP16 format, and stores them in Compressed Sparse Row (CSR) format for efficient computation. Besides, LLM-MQ models the bit-width assignment to each layer as an integer programming problem, and employs an efficient solver to solve it within a few seconds. Moveover, LLM-MQ designs a efficient CUDA kernel to integrate dequantization operators, thereby reducing memory access cost during computation.
@@ 5.2.2 Sparsification
The common focus of this line of work is the pruning criterion, including the weight importance and pruning ratio. Considering the huge parameter size of LLMs, improving the pruning efficiency is also crucial. One pruning criterion is to minimize the reconstruction loss of the model. SparseGPT [165] is a representative approach in this field. It follows the idea of OBS [219], which considers the impact of removing each weight on the network’s reconstruction loss. OBS iteratively decides a pruning mask to prune the weights and reconstructs the unpruned weights to compensate for the pruning loss. SparseGPT overcomes the efficiency bottleneck of OBS via the Optimal Partial Updates technique, and designs an adaptive mask selection technique based on the OBS reconstruction error. Prune and Tune [168] improves upon SparseGPT by fine-tuning the LLMs with minimal training steps during pruning. ISC [167] designs a novel pruning criterion by combining the saliency criteria in OBS [219] and OBD [220]. It further assigns non-uniform pruning ratios to each layer based on Hessian information. oBERT [171] and FastPruning [172] utilizes the second-order information of the loss function to decide the pruned weights. BESA [170] learns a differentiable binary mask via gradient descent of the reconstruction loss. The pruning ratio for each layer is sequentially decided by minimizing the reconstruction error. The other popular pruning criterion is magnitude-based. Wanda [166] proposes to use the element-wise product between the weight magnitude and the norm of input activation as the pruning criterion. RIA [173] jointly considers the weights and activations by using the metric of Relative Importance and Activations, which evaluates the importance of each weight element based on all its connected weights. In addition, RIA converts the unstructured sparsity pattern to a structured N:M sparsity pattern, which can enjoy the actual speed-up on NVIDIA GPUs. Additionally, OWL [169] focuses on deciding the pruning ratio of each layer. It assigns the pruning ratios to each layer based on its activation outlier ratios.
@@ 6 System-level Optimization
The system-level optimization for LLM inference primarily involves enhancing the model forward pass. Considering the computational graph of a LLM, there exist multiple operators, with attention and linear operators dominating most of the runtime. As mentioned in Sec. 2.3, system-level optimization primarily considers the distinctive characteristics of the attention operator and the decoding approach within LLM. In particular, to address the specific issues related to the decoding approach of LLMs, the linear operator requires special tiling designs, and speculative decoding methods are proposed to improve the utilization. Furthermore, in the context of online serving, requests come from multiple users. Therefore, beyond the optimizations discussed earlier, online serving faces challenges related to memory, batching and scheduling arising from asynchronous requests.
@@ 6.1 Inference Engine
The optimizations for inference engines are dedicated to accelerate the model forward process. Main operators and the computational graph in LLM inference are highly optimized. Besides, speculative decoding technique is proposed to accelerate the inference speed without performance degradation.
@@ 6.1.1 Graph and Operator Optimization
Runtime Profiling. Using HuggingFace [250] implementation, we profile the inference runtime with different models and context lengths. The profiling results in Fig. 15 demonstrate that attention operators and linear operators collectively dominate runtime, with their combined duration often exceeding 75% of the inference duration. Consequently, a significant portion of optimization efforts at the operator level is dedicated to enhancing the performance of the two operators. Furthermore, there are multiple operators occupying a small proportion of runtime, which fragments the operator execution timeline and increases the cost of kernel launch on the CPU side. To address this issue, at the computational graph level, current optimized inference engines implement highly fused operator.
@@
Recently, the application of the MoE FFN to enhance the model capability has become a trend in LLMs [12]. This model structure also puts forward new requirements for operator optimization. As shown in Fig. 15, in the Mixtral model with MoE FFN, the linear operator dominates the runtime due to the non-optimized FFN computation in the HuggingFace implementation. Besides, Mixtral’s adoption of the GQA attention structure decreases the attention operator’s runtime proportion, which further points out the urgent need to optimize the FFN layer. MegaBlocks [244] is the first to optimize the computation for MoE FFN layers. The work formulates the MoE FFN computation into block-sparse operations and proposes tailored GPU kernels for acceleration. However, MegaBlocks concentrates on the efficient training of the MoE models and hence ignores the characteristics of inference (e.g.,, the decoding approach). Existing frameworks are working hard to optimize the computations of the MoE FFN inference stage. The official repository of vLLM [51] integrates the fused kernels for MoE FFN in Triton [257], seamlessly removing the index overhead.
@@
The technique of kernel fusion proves effective with LLM inference, with all of the aforementioned benefits. FlashAttention [245] formulates the attention operator into one single kernel, removing the overhead of accessing the attention results. Based on the fact that the attention operator is memory-bounded, the reduction of memory access effectively transfers to runtime speed-up. ByteTransformer [247] and DeepSpeed [248] propose to fuse lightweight operators including residual adding, layernorm and activation functions, into the former linear operators to reduce the kernel launching overhead. As a result, those lightweight operators disappear in the timeline with nearly no extra latency. Moreover, kernel fusion is also adopted to enhance the utilization of LLM inference. The projections of Query, Key and Value matrices are originally three individual linear operations, and are fused into one linear operator to deploy on modern GPUs. Currently, the kernel fusion technique has been exploited in LLM inference practice, and highly optimized inference engines employ only a few fused kernels within the runtime. For example, in FlashDecoding++ [243] implementation, a transformer block integrates merely seven fused kernels. Leveraging the aforementioned operators and kernel fusion optimization, FlashDecoding++ achieves up to 4.86$\times$ speed-up over the HuggingFace implementation.
@@ 6.1.2 Speculative Decoding
Another line of studies focus on designing more effective draft construction strategies. Conventional approaches often yield single draft token sequences, posing challenges for passing verification. In response, Spectr [241] advocates generating multiple draft token sequences and employs a $k$-sequential draft selection technique to concurrently verify $k$ sequences. This method leverages speculative sampling, ensuring equivalence in output distributions. Similarly, SpecInfer [236] adopts a comparable approach. However, unlike Spectr, SpecInfer merges draft token sequences into a “token tree” and introduces a tree attention mechanism for validation. This strategy is called the ”token tree verifier”. Due to its efficacy, token tree verifier has been widely embraced in numerous speculative decoding algorithms [237, 235, 50, 240]. In addition to these efforts, Stage Speculative Decoding [237] and Cascade Speculative Drafting (CS Drafting) [238] propose accelerating draft construction by integrating speculative decoding directly into the token generation process.
@@
Tab. V provides a comparison of various speculative decoding methods, highlighting several key observations: (1) Eagle demonstrates exceptional performance, achieving a notable 3.47$\sim$3.72$\times$ end-to-end speed-up across multiple LLMs. To understand its success, a deeper analysis of Eagle reveals two key factors. Firstly, Eagle employs an auto-regressive approach for decoding draft tokens, leveraging information from previously generated tokens directly. Secondly, Eagle integrates rich features from previous tokens of both original LLMs and draft models to enhance the accuracy of next draft token generation. (2) The token tree verifier proves to be an effective technique in enhancing the performance of speculative decoding methods. (3) The end-to-end speed-up achieved by these methods is often lower than the acceptance rate. This difference arises due to the practical consideration that the generation cost associated with draft models cannot be overlooked.
@@ 6.2 Serving System
The optimizations for serving systems are dedicated to improve the efficiency in handling asynchronous requests. The memory management is optimized to hold more requests, and efficient batching and scheduling strategies are integrated to improve the system throughput. Besides, optimizations specific to distributed systems are proposed to exploit distributed computational resources.
@@ 6.2.3 Scheduling Strategy
In LLM serving, the job length of each request exhibits variability, and hence the order of executing requests significantly impacts the throughput of the serving system. The head-of-line blocking [269] happens when long requests are accorded priority. Specifically, memory usage grows rapidly in response to long requests, resulting in the impeding of subsequent requests when the system exhausts its memory capacity. The pioneering work ORCA [266] and open-source systems, including vLLM [51] and LightLLM [267], employ the simple first-come-first-serve (FCFS) principle to schedule requests. DeepSpeed-FastGen [268] gives priority to the decoding requests to enhance the performance. FastServe [269] proposes a preemptive scheduling strategy to optimize the head-of-line blocking problem, achieving low job completion time (JCT) in LLM serving. FastServe employs a multi-level feedback queue (MLFQ) to prioritize the requests with the shortest remaining time. Since the auto-regressive decoding approach poses unknown request lengths, FastServe predicts the length first and utilizes a skip-join fashion to find the proper priority for each request. Unlike previous work, VTC [270] discusses the fairness in LLM serving. VTC introduces a cost function based on token numbers to measure fairness among clients, and further proposes a fair scheduler to ensure fairness.

### v3

@@ 2.3 Efficiency Analysis
(Sec. 6.1), snode [Speculative Decoding, snode] [Offloading, snode] [Graph and Operator Optimization, snode] ] ] [Model-level Optimization (Sec. 5), mnode_1 [Model Compression
@@ 5.2.1 Quantization
In weight-only quantization, GPTQ [192] represents an early advancement in LLM quantization, building upon the traditional algorithm OBQ [221]. OBQ utilizes an optimal quantization order per row of the weight matrix, guided by the reconstruction error relative to the Hessian matrix of unquantized weights. After each quantization step, OBQ iteratively adjusts the unquantized weights to mitigate reconstruction errors. However, the frequent updating of the Hessian matrix during quantization escalates computational complexity. GPTQ streamlines this process by adopting a uniform left-to-right order for quantizing each row, thus circumventing the need for extensive Hessian matrix updates. This strategy substantially reduces computational demands by computing the Hessian matrix solely during the quantization of one row, then leveraging the computing results for subsequent rows, expediting the overall quantization procedure. LUT-GEMM [193] presents a novel dequantization method utilizing a Look-Up Table (LUT), aiming to accelerate the inference process of quantized LLMs by reducing the dequantization overhead. Additionally, it adopts a non-uniform quantization approach known as Binary-Coding Quantization (BCQ), which incorporates learnable quantization intervals. AWQ [194] observes that weight channels vary in importance for performance, particularly emphasizing those aligned with input channels exhibiting outliers in activations. To enhance the preservation of critical weight channels, AWQ utilizes a reparameterization method. This technique selects reparameterization coefficients via grid search to minimize reconstruction errors effectively. OWQ [195] observes the difficulty of quantizing weights associated with activation outliers. To address this challenge, OWQ employs a mixed-precision quantization strategy. This method identifies weak columns in the weight matrix and allocates higher precision to these specific weights, while quantizing the rest of the weights at a lower precision level. SpQR [196] introduces a methodology where weight outliers are identified and allocated higher precision during quantization, while the rest of the weights are quantized to 3 bits. SqueezeLLM [197] proposes to store the outliers in a full-precision sparse matrix, and apply non-uniform quantization to the remaining weights. The values for non-uniform quantization are determined based on quantization sensitivity, which contributes to improved performance of the quantized model. QuIP [198] introduces LDLQ, an optimal adaptive method for a quadratic proxy objective. The study reveals that ensuring incoherence between weight and Hessian matrices can enhance the effectiveness of LDLQ. QuIP utilizes LDLQ and achieves incoherence by employing random orthogonal matrix multiplication. FineQuant [199] utilizes a heuristic approach to determine the granularity of quantization per column, combining empirical insights gained from experiments to design a quantization scheme. QuantEase [200] builds upon GPTQ. When quantizing each layer, it proposes a method based on Coordinate Descent to compensate for the unquantized weights more precisely. Additionally, QuantEase can leverage quantized weights from GPTQ as an initialization and further refine the compensation process. LLM-MQ [201] protects the weight outliers with FP16 format, and stores them in Compressed Sparse Row (CSR) format for efficient computation. Besides, LLM-MQ models the bit-width assignment to each layer as an integer programming problem, and employs an efficient solver to solve it within a few seconds. Moveover, LLM-MQ designs a efficient CUDA kernel to integrate dequantization operators, thereby reducing memory access cost during computation. Inspired by the equivalent transformations used in the previous PTQ methods, AffineQuant [215] firstly introduces equivalent affine transformations in quantization, which extends the optimization scope and further reduces the quantization errors. Recently, many studies [198, 216, 218, 217] follows the computational invariance idea, by multiplying rotation matrices to the weight matrices and activation matrices. In this way, they can effectively eliminate the outliers in the weights and activations, thus help to quantize the LLMs. These studies use different rotation matrices. For example, QuaRot [217] applies randomize Hadamard transformations to the weights and activations. SpinQuant [218] finds the optimal rotation matrices by training on a small validation dataset.
@@ 5.2.2 Sparsification
The common focus of this line of work is the pruning criterion, including the weight importance and pruning ratio. Considering the huge parameter size of LLMs, improving the pruning efficiency is also crucial. One pruning criterion is to minimize the reconstruction loss of the model. SparseGPT [165] is a representative approach in this field. It follows the idea of OBS [225], which considers the impact of removing each weight on the network’s reconstruction loss. OBS iteratively decides a pruning mask to prune the weights and reconstructs the unpruned weights to compensate for the pruning loss. SparseGPT overcomes the efficiency bottleneck of OBS via the Optimal Partial Updates technique, and designs an adaptive mask selection technique based on the OBS reconstruction error. Prune and Tune [168] improves upon SparseGPT by fine-tuning the LLMs with minimal training steps during pruning. ISC [167] designs a novel pruning criterion by combining the saliency criteria in OBS [225] and OBD [226]. It further assigns non-uniform pruning ratios to each layer based on Hessian information. oBERT [171] and FastPruning [172] utilizes the second-order information of the loss function to decide the pruned weights. BESA [170] learns a differentiable binary mask via gradient descent of the reconstruction loss. The pruning ratio for each layer is sequentially decided by minimizing the reconstruction error. The other popular pruning criterion is magnitude-based. Wanda [166] proposes to use the element-wise product between the weight magnitude and the norm of input activation as the pruning criterion. RIA [173] jointly considers the weights and activations by using the metric of Relative Importance and Activations, which evaluates the importance of each weight element based on all its connected weights. In addition, RIA converts the unstructured sparsity pattern to a structured N:M sparsity pattern, which can enjoy the actual speed-up on NVIDIA GPUs. The recent study, Pruner-Zero [185], proposes to automatically identify the optimal pruning metric for LLMs, going beyond the hand-designed matrics. As a result, the optimal metric tailored for LLaMA and LLaMA-2 is $\boldsymbol{W}\odot\boldsymbol{W}\odot\sigma(\boldsymbol{G})$, where $\boldsymbol{W}$ and $\boldsymbol{G}$ represent the weights and gradients, and $\sigma(\cdot)$ scales a tensor to [0,1] using its mininum and maximum value. Additionally, OWL [169] focuses on deciding the pruning ratio of each layer. It assigns the pruning ratios to each layer based on its activation outlier ratios. $\O$[186] proposes a training-free appoarch to fine-tune the pruned LLMs. It builds upon the “pruning-and-growing” workflow adopted in Dynamic Sparse Training [227], which first prunes the model and then iteratively adjusts the network topology without training or weight update. $\O$T further designs the pruning and growing metrics tailored for LLMs.
@@ 6 System-level Optimization
The system-level optimization for LLM inference primarily involves enhancing the model forward pass. Considering the computational graph of an LLM, there exist multiple operators, with attention and linear operators dominating most of the runtime. As mentioned in Sec. 2.3, system-level optimization primarily considers the distinctive characteristics of the attention operator and the decoding approach within LLM. In particular, to address the specific issues related to the decoding approach of LLMs, the linear operator requires special tiling designs, and speculative decoding methods are proposed to improve the utilization. The substantial memory demand of LLMs leads to the offloading of parameters or KV cache to the CPU. Furthermore, in the context of online serving, requests come from multiple users. Therefore, beyond the optimizations discussed earlier, online serving faces challenges related to memory, batching, and scheduling arising from asynchronous requests.
@@ 6.1 Inference Engine
The optimizations for inference engines are dedicated to accelerating the model forward process. The main operators and the computational graph in LLM inference are highly optimized. Besides, the speculative decoding technique is proposed to accelerate the inference speed without performance degradation, and the offloading technique is introduced to mitigate the memory pressure.
@@ 6.1.1 Graph and Operator Optimization
Runtime Profiling. Using HuggingFace [260] implementation, we profile the inference runtime with different models and context lengths. The profiling results in Fig. 15 demonstrate that attention operators and linear operators collectively dominate runtime, with their combined duration often exceeding 75% of the inference duration. Consequently, a significant portion of optimization efforts at the operator level is dedicated to enhancing the performance of the two operators. Furthermore, there are multiple operators occupying a small proportion of runtime, which fragments the operator execution timeline and increases the cost of kernel launch on the CPU side. To address this issue, at the computational graph level, current optimized inference engines implement highly fused operators.
@@
Recently, the application of the MoE FFN to enhance the model capability has become a trend in LLMs [12]. This model structure also puts forward new requirements for operator optimization. As shown in Fig. 15, in the Mixtral model with MoE FFN, the linear operator dominates the runtime due to the non-optimized FFN computation in the HuggingFace implementation. Besides, Mixtral’s adoption of the GQA attention structure decreases the attention operator’s runtime proportion, which further points out the urgent need to optimize the FFN layer. MegaBlocks [254] is the first to optimize the computation for MoE FFN layers. The work formulates the MoE FFN computation into block-sparse operations and proposes tailored GPU kernels for acceleration. However, MegaBlocks concentrates on the efficient training of the MoE models and hence ignores the characteristics of inference (e.g., the decoding approach). Existing frameworks are working hard to optimize the computations of the MoE FFN inference stage. The official repository of vLLM [51] integrates the fused kernels for MoE FFN in Triton [267], seamlessly removing the index overhead.
@@
The technique of kernel fusion proves effective with LLM inference, with all of the aforementioned benefits. FlashAttention [255] formulates the attention operator into one single kernel, removing the overhead of accessing the attention results. Based on the fact that the attention operator is memory-bounded, the reduction of memory access effectively transfers to runtime speed-up. ByteTransformer [257] and DeepSpeed [258] propose to fuse lightweight operators including residual adding, layernorm, and activation functions, into the former linear operators to reduce the kernel launching overhead. As a result, those lightweight operators disappear in the timeline with nearly no extra latency. Moreover, kernel fusion is also adopted to enhance the utilization of LLM inference. The projections of Query, Key, and Value matrices are originally three individual linear operations, and are fused into one linear operator to deploy on modern GPUs. Currently, the kernel fusion technique has been exploited in LLM inference practice, and highly optimized inference engines employ only a few fused kernels within the runtime. For example, in FlashDecoding++ [253] implementation, a transformer block integrates merely seven fused kernels. Leveraging the aforementioned operators and kernel fusion optimization, FlashDecoding++ achieves up to 4.86$\times$ speed-up over the HuggingFace implementation.
@@ 6.1.2 Speculative Decoding
Another line of studies focuses on designing more effective draft construction strategies. Conventional approaches often yield single draft token sequences, posing challenges for passing verification. In response, Spectr [248] advocates generating multiple draft token sequences and employs a $k$-sequential draft selection technique to concurrently verify $k$ sequences. This method leverages speculative sampling, ensuring equivalence in output distributions. Similarly, SpecInfer [243] adopts a comparable approach. However, unlike Spectr, SpecInfer merges draft token sequences into a “token tree” and introduces a tree attention mechanism for validation. This strategy is called the ”token tree verifier”. Due to its efficacy, token tree verifier has been widely embraced in numerous speculative decoding algorithms [244, 242, 50, 247]. In addition to these efforts, Stage Speculative Decoding [244] and Cascade Speculative Drafting (CS Drafting) [245] propose accelerating draft construction by integrating speculative decoding directly into the token generation process.
@@
Tab. V provides a comparison of various speculative decoding methods, highlighting several key observations: (1) Eagle demonstrates exceptional performance, achieving a notable 3.47$\sim$3.72$\times$ end-to-end speed-up across multiple LLMs. To understand its success, a deeper analysis of Eagle reveals two key factors. Firstly, Eagle employs an auto-regressive approach for decoding draft tokens, leveraging information from previously generated tokens directly. Secondly, Eagle integrates rich features from previous tokens of both original LLMs and draft models to enhance the accuracy of the next draft token generation. (2) The token tree verifier proves to be an effective technique in enhancing the performance of speculative decoding methods. (3) The end-to-end speed-up achieved by these methods is often lower than the acceptance rate. This difference arises due to the practical consideration that the generation cost associated with draft models cannot be overlooked.
6.1.3 Offloading
Current research investigates the potential of offloading to accommodate the substantial memory demand of LLMs (see Sec. 2.3) in resource-constrained environments. The essence of offloading is to offload part of the storage from the GPU to the CPU when it is free of use. Intuitively, the focus of such kind of research lies in hiding the expensive data movement latency between the GPU and the CPU. FlexGen [203] enables the offloading of weights, activations, and the KV cache, and further formulates a graph traversal problem for offloading to maximize the throughput. The data loading of the next batch and the data storing of the previous batch can be overlapped with the computation of the current batch. Another work llama.cpp [250] also assigns computational tasks to the CPU, mitigating the data transfer overhead at the cost of computing with the low-powered CPU. Powerinfer [251] exploits the sparsity in activations using ReLU [271] in LLMs, and divides the activations into subsets of cold and hot neurons representing the frequency of computation. The cold neurons are offloaded to the CPU for both storage and computation in Powerinfer. Leveraging adaptive predictors and sparse operators, Powerinfer significantly improves the computational efficiency with offloading. FastDecode [252] proposes to offload the storage and the computation of the entire attention operator to the CPU. Since the attention operation is computed on the CPU, the data movement of KV cache is reduced to merely some activations. The number of CPUs is selected to match the workload latency on GPUs so that the bubbles in the heterogeneous pipeline are mitigated.
@@ 6.2 Serving System
The optimizations for serving systemworks are dedicated to improving the efficiency in handling asynchronous requests. The memory management is optimized to hold more requests, and efficient batching and scheduling strategies are integrated to enhance the system throughput. Besides, optimizations specific to distributed systems are proposed to exploit distributed computational resources.
@@ 6.2.3 Scheduling Strategy
In LLM serving, the job length of each request exhibits variability, and hence the order of executing requests significantly impacts the throughput of the serving system. The head-of-line blocking [280] happens when long requests are accorded priority. Specifically, memory usage grows rapidly in response to long requests, impeding subsequent requests when the system exhausts its memory capacity. The pioneering work ORCA [277] and open-source systems, including vLLM [51] and LightLLM [278], employ the simple first-come-first-serve (FCFS) principle to schedule requests. DeepSpeed-FastGen [279] gives priority to the decoding requests to enhance the performance. FastServe [280] proposes a preemptive scheduling strategy to optimize the head-of-line blocking problem, achieving low job completion time (JCT) in LLM serving. FastServe employs a multi-level feedback queue (MLFQ) to prioritize the requests with the shortest remaining time. Since the auto-regressive decoding approach poses unknown request lengths, FastServe predicts the length first and utilizes a skip-join fashion to find the proper priority for each request. Unlike previous work, VTC [281] discusses the fairness in LLM serving. VTC introduces a cost function based on token numbers to measure fairness among clients, and further proposes a fair scheduler to ensure fairness.